how many calories are needed to raise the temperature of 21.0 g of silver from 22 degrees Celsius to 48 degrees Celcius
1
Expert's answer
2020-05-03T17:18:29-0400

Let find the energy, that need to heat silver:

"Q=c\times m\times \Delta T" ,

where:

"\Delta T=T_{2}-T_{1}=26K, m=21g=21\times10^{-3} kg."

c is a constant and for silver:

"c=250\dfrac{J}{kg\times K}"

"Q=136.5 J"

We know that:

"1c(calorie)=4.184J(Joule)"

or:

"1J(Joule)=0.239c(calorie)"

So to raise the temperature of silver we need:

"0.239\times 136.5=32.6235 c(calories)"